OpenStudy (anonymous):

yo help me out Is the line through points P(–8, –10) and Q(–5, –12) perpendicular to the line through points R(9, –6) and S(17, –5)? Explain

OpenStudy (anonymous):

find the slopes of both lines and see if they are opposite reciprocals of each other

OpenStudy (anonymous):

i would if i knew, can you please give me the answer ill give u a medal

OpenStudy (anonymous):

to find the slope you use this formula\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

so for the first two points you get [-12-(-10)]/[-5-(-8)] =-2/3

OpenStudy (anonymous):

do the same for 2nd line and compare

OpenStudy (anonymous):

alrght let me try